OpenStudy (anonymous):

a. Solve this rational equation with steps 2x+12/x-2=x/x-2+5

ganeshie8 (ganeshie8):

\[\large 2x + \dfrac{12}{x-2} = \dfrac{x}{x-2} + 5\] like this ?

OpenStudy (anonymous):

yes i have no idea how to solve this

ganeshie8 (ganeshie8):

just to be 100% sure, could you take a screenshot of the equation and attach if psble ?

ganeshie8 (ganeshie8):

thnks :) since (x-2) is in the bottom, lets get rid of it by multiplying (x-2) through out the equation

ganeshie8 (ganeshie8):

\[\large 2x\color{gray}{(x-2)} + \dfrac{12}{x-2}\color{gray}{(x-2)} = \dfrac{x}{x-2}\color{gray}{(x-2)} + 5\color{gray}{(x-2)} \]

ganeshie8 (ganeshie8):

canceling you get : \[\large 2x\color{gray}{(x-2)} + 12 =x + 5\color{gray}{(x-2)} \]

ganeshie8 (ganeshie8):

distribute and simplify as much as possible

OpenStudy (anonymous):

Thank you love :) <3

ganeshie8 (ganeshie8):

np :) are you sure you can manage the rest ?

OpenStudy (anonymous):

yes i can sweetie :) thank you so much

ganeshie8 (ganeshie8):

sounds good, you will get a quadratic which you might be able to solve by factoring and make sure to discard "2" if you get it in solutions because x=2 makes the given equation go crazy (why ?)